PartnerObserving Log for 2007 Jan 20/21

Observers: Rick Pogge & Xiaohui Fan
Inst. Support: Dave Thompson
TOs: Aaron Ceranski & David Gonzalez

1800 UTC Jan 20: About a foot or so of snow on the summit, though high winds swept it mostly clear of the LBT surroundings. All the buildings and trees are rimed with snow and ice. Very pretty, but not that promising for tonight judging from the satellite. Fog enshrouded us shortly after taking some pretty pictures when the sun peeked out.

  • LBT the morning after (2007 Jan 20):
    SnowyLBT070120.jpg

All dates below are UTC 2007 Jan 21:

0040 - Foggy and below freezing all afternoon until sunset, when we got a peek at the Sun briefly before the clouds rolled back in. Humidity >95% at sunset.

Building rotation is back, so we should be in business if it clears, or at least we'll be able to work through the OBs to familiarize ourselves with the system.

0142 - Fired up LBCB and commenced taking biases (25bias.ob) at 0148. Dome still closed due to icy fog.

0225 - Turned on dome lights, created new OB for a 1/2/4/8sec ramp in Bessel-V, just to exercise making new OBs on the fly and executing them. 8s saturated CCDs

0235 - Second test ramp, 1/2/4/6/0.5, but last 0.5s exposure was really 1 second! (1.24 in header). Counts in the image are consistent with the 1s exposure. Noted as IssueTrak #440

0245 - We can see Orion, but humidity is still 94%, Wind 3-5m/s, Temp -11C

0255 - Ran a ramp 1/2/3/4/5s exposure Bessel-V with dome lights on. First run the 3s exposure in the sequence never made it into the respository (where it would be between lbcb.20070121.025857 and .030016). Ran the test OB a second time starting at 0304 and all 5 images were acquired and stored in the repository. The missing image is associated with a "too few downloaded pixels" error at 03:00:10.404 in the log (as per IssueTrak #375).

0350 - Clear, but 94% humidity and some blowing bits of junk. We'll do some mock observing just to exercise the telescope and OBs and work through the sequences. Orion looks really nice, but the satellite shows it's just a sucker hole, with more clouds coming. TOs firing up the LBT.

0425 - mock observation of Sigma Ori Field 1, started with mock WF check, then executing sdt_muzerolle_1.ob

0455 - continuing mock observation with sdt_muzerolle_2.ob, all systems are working smoothly.

0525 - mock WF check, then mock observation of sdt_muzerolle_3.ob

0552 - mock observation of sdt_muzerolle_4.ob, successfully "completing" Sigma Ori Field1 in both rotations. Humidity still around 95%. Final observation completed at 0620. Total execution time about 2h, as predicted. Everything worked well except the weather, no glitches or other problems with LBC or the telescope.

0900 - Instrument systems turned off and deauthorized. Night is pretty much a loss, despite the teasing starlight behind the high humidity.

-- RichardPogge - 20/21 Jan 2007
Topic revision: r11 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback